Use dynamic programming algorithm to compute the binary coefficient C(7,5). Show all steps!
Answer #1

in DP for binary coefficient ,

Recurrence relaton is:

C(n,k) = C(n-1,k) + C(n-1,k-1) for n > k > 0

                      C(n,0) = 1,   C(n,n) = 1 for n >= 0

we will make table, value of C(7,5) will be found from table

C(7,5) = C(6,5)+C(6,4) because [C(n,k) = C(n-1,k) + C(n-1,k-1) ]

c(column-wise)

n(row-wise)

0 1 2 3 4 5
0
1
2
3
4
5
6
7

Step 1:   C(n,0) = 1,   C(n,n) = 1 for n >= 0

c(column-wise)

n(row-wise)

0 1 2 3 4 5
0 1
1 1 1
2 1 1
3 1 1
4 1 1
5 1 1
6 1
7 1

Step 2: C(n,k) = C(n-1,k) + C(n-1,k-1) for n > k > 0

c(column-wise)

n(row-wise)

0 1 2 3 4 5
0 1
1 1 1
2 1 2 1
3 1 3 3 1
4 1 4 6 4 1
5 1 5 10 10 5 1
6 1 6 15 20 15 6
7 1 7 21 35 35 21

C(7,5) = C(6,5)+C(6,4) = 6+15 =21